OpenStudy (anonymous):

A quadratic equation is shown below: x2 - 8x + 13 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? Subtract 5 from both sides of the equation Subtract 3 from both sides of the equation Add 5 to both sides of the equation Add 3 to both sides of the equation

myininaya (myininaya):

do you know how to complete the square? let's practice do you think x^2-2x+1 can be written as a square?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

i was thnking it was D?

myininaya (myininaya):

can you write is as a square for me?

OpenStudy (anonymous):

(X-1)^2

myininaya (myininaya):

okay so if you do that then we have x^2-8x+13+3=0+3 x^2-8x+16=3

myininaya (myininaya):

so are you think x^2-8x+16 can be written as square?

myininaya (myininaya):

and you are right about my example

OpenStudy (anonymous):

Yes

myininaya (myininaya):

x^2-8x+16 can be written as (x- ? )^2

myininaya (myininaya):

what is ?

OpenStudy (anonymous):

(x-4)^2

myininaya (myininaya):

yeah

myininaya (myininaya):

you did this all by yourself @kristinak

OpenStudy (anonymous):

yeah but i still dont know the answer

myininaya (myininaya):

you are saying we have (x-4)^2=3

myininaya (myininaya):

4 and 3 are totally integers

OpenStudy (anonymous):

so its b?

myininaya (myininaya):

:(

myininaya (myininaya):

Can you tell me where you are confused on your first choice?

OpenStudy (anonymous):

i thought you said i was wrong when i said it was D?

myininaya (myininaya):

never said that

myininaya (myininaya):

i simply was working through what D would have given us

myininaya (myininaya):

do you not think 4 and 3 are integers?

OpenStudy (anonymous):

No they are

OpenStudy (anonymous):

thank you so much, sorry i was confused

myininaya (myininaya):

its okay

OpenStudy (anonymous):

can you help me with another?

myininaya (myininaya):

alright

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

its the screenshot

myininaya (myininaya):

can you find the vertex for the first one?

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

(70,55)

OpenStudy (anonymous):

am i right?

myininaya (myininaya):

beautiful!

myininaya (myininaya):

now the vertex is very important for consider the min or even the max for a parabola that is why i asked for the other graph i think a good way to approximate the vertex is just plot those points given

myininaya (myininaya):

i want you to speculate on what you think the vertex is

myininaya (myininaya):

now it helped me to actually plot the points so i encourage you to do so

myininaya (myininaya):

tell me what you think looks like the middle point of that graph

OpenStudy (anonymous):

(1,899.50)

myininaya (myininaya):

that is very great

myininaya (myininaya):

now the y values for each of these points is the amount of min production cost from each graph

myininaya (myininaya):

aka lowest y value for each graph

myininaya (myininaya):

the question ask you to compare the y values basically it ask which one is greater

OpenStudy (anonymous):

company 2?

myininaya (myininaya):

right because 899.5>55 so company 2 has a bigger min production cost

OpenStudy (anonymous):

thank you!! can you help me with like 2 or 3 more?

myininaya (myininaya):

honestly i cannot i have to leave i suggest posting a new question

OpenStudy (anonymous):

oh okay thank you

myininaya (myininaya):

np have a great night